OpenStudy (prominer10000):

PLEASE HELP! WILL MEDAL AND FAN! What is the point-slope form of the equation of a line that passes through the point (6, -8) and has a slope of -2? y+8=2(x-6) y+8=-2(x-6) y-6=-2(x-8) y-8=-2(x+6) I'm thinking D?

OpenStudy (will.h):

same as before we will have to use the following form y - y1 = m(x - x1) where m is the slope and we have the point (6,-8) where y1 is -8 and x1 is 6 and we know that m is -2 so substitute y - (-8) = -2(x - 6) simplify the left side y + 8 = -2(x - 6)